Someone please help me, I am completely clueless about linear equations.
1. Solve the system of linear equations by graphing: x+2y=5, x-y=2
A. (1,3)
B. (3,1)
C. (7,5)
D. (5,7)
2. Solve the system of linear equations by graphing: 4x-y=6, 3x+y=1
A. (-2,1)
B. (1,-2)
C. (5,14)
D. (2,2)
3. Solve the system of equations by graphing: y=x^2-4, y=x-2
A. (2,0) and (-1,-3)
B. (2,0) and (1,-3)
C. (3,5) and (-1,-3)
D. (3,5) and (1,-3)
Thank you so much for all your help!!!!!!

Answer by rfer(16322)

we will solve number one by substitution
x-y=2, same as x=y+2
x+2y=5
substitute the first equation into the second
y+2+2y=5
subtract 2
3y=3
y=1
-------------------------------
work the first equation using y=1
x-1=2
add 1
x=3
ordered pair
(3,1)point of intersection of the two equations
------------------------
use y=mx+b for the other points
example
x+2y=5
2y=-x+5
y=-1/2+5/2
m=slope
b=y int
Give the others a try.
Bob